# Counterfactual Explanations with Probabilistic Guarantees on their Robustness to Model Change

### v1

@@ 4.1 Experimental setup
We conducted experiments [^3] To ensure full reproducibility, we open-source the code (see App. A.2) on four datasets: HELOC, Wine, Diabetes, and Breast Cancer, which are commonly used in related studies. Each dataset varies in size and number of attributes, but all are numerical to meet the requirements of the methods used later. Brief characteristics of these datasets are presented in Tab. 1.
Table 1: Datasets characteristics.
Dataset Examples Features Target
Wine 6 497 11 binary [^4] The discrete target feature, binarized by $x>5$
Diabetes 768 8 binary
HELOC 2 502 23 binary
Breast Cancer 569 30 binary
We conducted experiments using two types of models: neural networks (Goodfellow et al., 2016) and gradient boosted trees (LightGBM (Ke et al., 2017)), with the results of the latter moved entirely to App. H. The neural network architecture comprised three layers, each consisting of 128 neurons and ReLU (Nair & Hinton, 2010) activation functions. The training procedure utilized a binary cross-entropy loss function and Adam optimizer (Kingma & Ba, 2015) with 1e-3 learning rate. Early stopping with patience of 5 epochs on validation loss was applied. The models were trained for a maximum of 100 epochs, with data batches of size 128555The comprehensive list of hyperparameters is in App. F.4. As base CFE generation methods, we used two popular algorithms: Dice  (Mothilal et al., 2020) and GrowingSpheres (Laugel et al., 2018) which were briefly introduced in Sec. 2.
@@
  • Seed – both the data and model parameters remain unchanged. The difference in models arises from using different initial random states, which impacts the decision boundary (Fellicious et al., 2020).
@@
In all experiments, we set the confidence level to $\alpha=0.95$. In the first experiment (Sec. 4.2), we opted to utilize credible intervals to demonstrate that both lower and upper bounds on the robustness hold in practice. Later, in the comparative analysis (Sec. 4.4) we only employed the lower bound, following the exact definition from Def. 4. Since $\alpha=0.95$ was used for credible intervals, the lower bound was set to $0.975$ for consistency (since it only uses the left tail). With this confidence level, we used the introduced formula (Eq. 7) to select the number of estimators ($k=32$) required to achieve the highest $\delta$ investigated in the experiments (i.e. $\delta=0.9$).
@@ 4.2 Validation of BetaRCE theoretical framework
In the first experiment, we validated whether the theoretical framework, from which we derived BetaRCE, holds in considered experimental settings. To achieve this, we opted not only to use the lower bound from our $(\delta,\alpha)$-robustness definition (Def. 4), but also calculate the upper bound to extract more information from the method. In Fig. 2 we present the results for two datasets, Diabetes and Wine. The base CFEs were generated using GrowingSpheres. Plots for other datasets and Dice serving as a baseline CFE generation method are available in App. G.1.
@@
Only for the Diabetes dataset and Seed scenario, we notice that the empirical robustness at $\delta=0.9$ is slightly lower than the estimated average lower bound (however still above the $\delta$). We attribute this discrepancy to the $95\%$ confidence level of the method, which acknowledges that the estimate may occasionally be incorrect.
@@ 4.3 Hyperparameter sensitivity analysis
In this section, we present the experiments assessing the impact of using different confidence values $\alpha$ and different numbers of estimators $k$. Selected results for varying $\alpha$ and $k$ are shown in Fig 3 and Fig. 4, respectively [^6] This analysis is extended in App. G.2 and G.3.
The results for different $\alpha$ values indicate that as $\alpha$ increases, the empirical robustness tends to move further away from the lower bound. This suggests that higher $\alpha$ values yield more confident estimates with a lower likelihood of violating the lower bound.
@@
Figure 3: The average empirical robustness computed at various confidence ($\alpha$) and robustness ($\delta$) levels on the HELOC dataset, with $k$ set to $32$.
Figure 4: The impact of the number of $k$ estimators on the width of $95\%$ credible interval, computed on the HELOC dataset.
@@ 4.4 Comparative study with other methods
Table 2: Comparative study results. RobX and BetaRCE were using GrowingSpheres as the base counterfactual explainer. Parameters used in a given method are listed next to this method’s name; for RobX these are $\tau$ and variance, while for BetaRCE $\delta$ and $\alpha$. The values in each cell represent the mean $\pm$ standard error. The column Type sorts the methods by categories. The abbreviations Btsr and Arch used next to BetaRCE in the Type column stand for Bootstrap and Architecture, respectively.
@@
BetaRCE Arch BetaRCE (0.8,0.975) 0.488 $\pm$ 0.013 0.870 $\pm$ 0.013 $0.382\pm 0.005$ 0.372 $\pm$ 0.004 0.966 $\pm$ 0.004 - -
BetaRCE (0.9,0.975) 0.607 $\pm$ 0.014 0.953 $\pm$ 0.013 $0.420\pm 0.006$ 0.378 $\pm$ 0.004 0.975 $\pm$ 0.004 - -
BetaRCE Btsr BetaRCE (0.8,0.975) 0.445 $\pm$ 0.006 0.840 $\pm$ 0.008 $0.359\pm 0.003$ 0.359 $\pm$ 0.002 - 0.903 $\pm$ 0.006 -
BetaRCE (0.9,0.975) 0.583 $\pm$ 0.007 0.949 $\pm$ 0.008 $0.407\pm 0.003$ 0.369 $\pm$ 0.002 - 0.937 $\pm$ 0.005 -
BetaRCE Seed BetaRCE (0.8,0.975) 0.247 $\pm$ 0.006 0.813 $\pm$ 0.008 $0.346\pm 0.003$ 0.350 $\pm$ 0.002 - - 0.871 $\pm$ 0.006
BetaRCE (0.9,0.975) 0.315 $\pm$ 0.006 0.862 $\pm$ 0.008 $0.367\pm 0.003$ 0.353 $\pm$ 0.002 - - 0.902 $\pm$ 0.006
@@
BetaRCE Arch BetaRCE (0.8,0.975) 1.538 $\pm$ 0.049 2.912 $\pm$ 0.053 $0.749\pm 0.014$ 0.802 $\pm$ 0.011 0.904 $\pm$ 0.006 - -
BetaRCE (0.9,0.975) 1.697 $\pm$ 0.031 2.927 $\pm$ 0.036 $0.753\pm 0.009$ 0.783 $\pm$ 0.007 0.935 $\pm$ 0.005 - -
BetaRCE Btsr BetaRCE (0.8,0.975) 2.288 $\pm$ 0.041 3.451 $\pm$ 0.044 $0.889\pm 0.011$ 0.859 $\pm$ 0.008 - 0.833 $\pm$ 0.007 -
BetaRCE (0.9,0.975) 3.547 $\pm$ 0.071 4.501 $\pm$ 0.073 $1.156\pm 0.019$ 1.044 $\pm$ 0.015 - 0.880 $\pm$ 0.006 -
BetaRCE Seed BetaRCE (0.8,0.975) 1.420 $\pm$ 0.021 2.526 $\pm$ 0.028 $0.653\pm 0.007$ 0.726 $\pm$ 0.004 - - 0.826 $\pm$ 0.007
BetaRCE (0.9,0.975) 1.927 $\pm$ 0.030 2.906 $\pm$ 0.035 $0.750\pm 0.009$ 0.776 $\pm$ 0.006 - - 0.902 $\pm$ 0.006
In the final experiment, we empirically compare the performance of BetaRCE with several baselines, each representing a different type of method and aimed at achieving different CFE properties (see Sec. 2). The first type of baselines includes standard CFE generation methods: Dice  (Mothilal et al., 2020), FACE  (Poyiadzi et al., 2020), and GrowingSpheres  (Laugel et al., 2018), which do not claim to guarantee any robustness to model change. The next type are end-to-end CFE generation methods, explicitly generating robust CFEs: RBR  (Nguyen et al., 2022) and ROAR  (Upadhyay et al., 2021). Finally, we also include RobX (Dutta et al., 2022); a method most similar to ours, as it also operates on top of a base CFE generation method in a post-hoc fashion. The above-mentioned robust baselines aim to increase overall robustness of CFEs by moving the CFE to a "safer" region in the feature space, requiring manual selection of hyperparameters in a dataset-specific manner. To ensure a fair comparison, for RBR and ROAR , we performed a hyperparameter search to find the most promising settings. In the case of RobX, we chose four combinations of the most relevant hyperparameters, $\tau$ and variance, where $\tau$ values were selected using the histogram technique described by the authors, and variances were chosen to vary significantly (specifically $0.1$ and $0.01$). The comprehensive description of the hyperparameter selection process is described in App. F.3. Note, that this comparison is not straightforward since each method is built on different premises and has multiple parameters to tune, making comparison to our probabilistic goals challenging.
We compare the above-listed baselines to two variants of our method ($\delta=0.8$, $\delta=0.9$) on several quality measures: Empirical robustness already defined in Eq. 8. Proximity: The distance between the counterfactual and the original example, $d(x^{cf},x^{orig})$. Notation $L_{1}$ and $L_{2}$ corresponds to Manhattan and Euclidean distance, respectively. Plausibility: The average distance to $n$ closest neighbors of the CFE in the training set, $\frac{1}{n}\sum_{i=1}^{n}d(x^{cf},x_{i})$. Distance to Base: A metric calculated only for the post-hoc methods, measuring the $L_{1}$ distance between the robust CFE and the base one: $d(x^{rcf},x^{bcf})$.
@@
The detailed results for Diabetes and HELOC datasets are presented in Tab. 2 and the results for the remaining datasets are in App. H. The results indicate that BetaRCE is outperforming RobX in terms of the Distance to Base metric. Moreover, both the Proximity and Plausibility metrics are either preserved or slightly improved, validating our goal of not deteriorating the properties of the base counterfactual. This suggests that BetaRCE is effective at finding CFEs with a user-defined robustness level that are not significantly different from the base counterfactual, preserving its original properties and aligning with one of our primary motivations.
@@
Regarding Empirical Robustness, we observe that RobX often achieves a higher score than BetaRCE. However, this comes at a cost, as indicated by the fact that the Distance to Base metric for RobX is frequently worse than that of BetaRCE. In some cases, RobX with well-selected hyperparameters achieves nearly perfect robustness but significantly deteriorates all other metrics (e.g. more than doubles the Distance to Base in comparison to BetaRCE ). Recall that the goal of BetaRCE is to have robustness greater than a specified lower bound while preserving as much of the base counterfactual as possible, and these goals have been achieved. The only setting where the lower bound is violated by the computed Empirical Robustness is the Bootstrap-HELOC combination for BetaRCE at $\delta=0.9$, where that value is slightly lower than the expected lower bound. However, the difference is negligible and we attribute it to BetaRCE’s confidence level parameter.
@@
We also conducted additional experiments to verify the robustness of BetaRCE to a misdefined space of admissible model changes. The details are presented in App. G.4. In summary, although the theoretical probabilistic bounds do not hold for out-of-distribution model changes, they are often practically satisfied by BetaRCE . The results reveal good transferability of CFEs generated for Architecture and Bootstrap scenarios to Seed, as well as Seed CFEs to Architecture. The transfer to the Bootstrap scenario seems to be more difficult, sometimes resulting in a drop of Empirical robustness as high as 4 p.p.
@@ 5 Final Remarks
The experiments have confirmed that the introduced probabilistic bounds, estimated in practice with a set of models, hold for the Empirical Robustness, as demonstrated across three types of model changes. Notably, counterfactuals generated by BetaRCE not only have improved robustness, but are also closer to the base CFEs and retain their properties better than those generated by existing methods, while achieving a similar level of robustness.
There are several possible avenues for future research. First, our approach currently uses a fairly simple bootstrap estimation method of the second-order probability distribution of robustness. Despite the fact that our experiments confirmed the high utility of this estimation approach, the use of more advanced estimation methods could lead to interesting results. For example, one could investigate the use of auxiliary models to predict beta distribution parameters instead of querying the set of estimators. Additionally, the feasibility of using different and faster optimization algorithms to speed up the optimization process could be investigated.
@@ Appendix A Reproducibility
A.1 Datasets
Wine and Breast Cancer datasets were obtained from https://archive.ics.uci.edu/, while the Diabetes dataset was sourced from https://www.kaggle.com/datasets/mathchi/diabetes-data-set, and the HELOC dataset from https://community.fico.com/s/explainable-machine-learning-challenge?tabset-158d9=d157e.
A.2 Code Availability
To ensure reproducibility and enable further experimentation with BetaRCE, we make the source code publicly available on GitHub (for the review period we anonymized the repository): https://anonymous.4open.science/r/KDD-1115.

### v3

@@ 4.1 Experimental setup
We conducted experiments on six datasets: HELOC, Wine, Diabetes, Breast Cancer, Car evaluation, and Rice – all commonly used in related studies. Each dataset varies in size and number of attributes, but all are numerical to meet the requirements of the methods used later. Brief characteristics of these datasets are presented in App. E.2. For reproducibility purposes, we open-source our code [^3] https://github.com/istepka/betarce.
We conducted experiments with three types of models: neural networks [32], gradient boosted trees (LightGBM [33]) and logistic regression [34]. The results for the latter two are provided in App. K. The neural network architecture comprised three layers, each consisting of 128 neurons and ReLU [35] activation functions. The training procedure utilized a binary cross-entropy loss function, Adam optimizer [36] with 1e-3 learning rate, and early stopping with patience of 5 epochs. The models were trained for a maximum of 100 epochs, with data batches of size 128 The comprehensive list of hyperparameters can be found in App. E.4. As base CFE generation methods, we used two popular algorithms: Dice  [18] and GrowingSpheres [16] which were briefly introduced in Sec. 2.
@@
  • Seed – both the data and model parameters remain unchanged. The difference in models arises from using different initial random states, which impacts the decision boundary [37].
@@
In all experiments, we set the confidence level to $\alpha=0.9$. In the first experiment (Sec. 4.2), we opted to utilize credible intervals to demonstrate that both lower and upper bounds on the robustness hold in practice. Later, in the comparative analysis (Sec. 4.4) we only employed the lower bound, following the exact definition from Def. 4. Since $\alpha=0.9$ was used for credible intervals, the lower bound was set to $0.95$ for consistency (as it only uses the left tail). With this confidence level, we used the introduced formula (Eq. 7) to select the number of estimators ($k=32$) required to achieve the highest $\delta$ (i.e., $\delta=0.9$) investigated in the experiments.
@@
Figure 2: The average Empirical Robustness of counterfactuals generated by BetaRCE at $\alpha=90\%$, with GrowingSpheres generating base CFEs. Red and green dashed lines are average lower and upper bounds of estimated $\alpha$ credible intervals (CI). The shaded areas in the back represent standard deviations. Horizontal yellow and gray lines show the base robustness of CFEs obtained with Dice and GrowingSpheres without BetaRCE applied. The plots with all four datasets and two BetaRCE base CFE generation methods (Dice and GrowingSpheres ) are available in App. F.1.
@@ 4.2 Validation of BetaRCE theoretical framework
In the first experiment, we validated whether the theoretical framework, from which we derived BetaRCE, holds in considered experimental settings. To achieve this, we opted not only to use the lower bound from our $(\delta,\alpha)$-robustness definition (Def. 4), but also calculate the upper bound to extract more information from the method. In Fig. 2 we present the results for two datasets, Diabetes and Wine. The base CFEs were generated using GrowingSpheres.
@@
Only for the Diabetes dataset and Seed scenario, we notice that the empirical robustness at $\delta=0.9$ is slightly lower than the estimated average lower bound (however still above the $\delta$). We attribute this discrepancy to the $90\%$ confidence level of the method, which acknowledges that the estimate may occasionally be incorrect.
@@ 4.3 Hyperparameter sensitivity analysis
In this section, we present the experiments assessing the impact of using different confidence values $\alpha$ and different numbers of estimators $k$. Selected results for varying $\alpha$ and $k$ are shown in Fig 3. The hyperparameter sensitivity analysis is extended in App. F.2 and F.3.
The results for different $\alpha$ values indicate that as it increases, the empirical robustness tends to move further away from the lower bound. This suggests that higher $\alpha$ values yield more confident estimates with a lower likelihood of violating the lower bound.
@@
Figure 3: Left: The average empirical robustness computed at various confidence ($\alpha$) and robustness ($\delta$) levels on the HELOC dataset, with $k$ set to $32$. Right: The impact of the number of $k$ estimators on the width of $90\%$ credible interval, computed on the HELOC dataset.
@@ 4.4 Comparative study with other methods
In the final experiment, we empirically compare the performance of BetaRCE with several baselines, each representing a different type of method and aimed at achieving different CFE properties (see Sec. 2). The first type of baselines includes standard CFE generation methods: Dice  [18], FACE  [17], and GrowingSpheres  [16], which do not claim to guarantee any robustness to model change. The next type are end-to-end CFE generation methods, explicitly generating robust CFEs: RBR  [25] and ROAR  [10]. Finally, we also include RobX [11]; a method most similar to ours, as it also operates on top of a base CFE generation method in a post-hoc fashion. The above-mentioned robust baselines aim to increase overall robustness of CFEs by moving the CFE to a more robust region in the feature space. Notably, they require manual selection of hyperparameters in a dataset-specific manner. Therefore, to ensure a fair comparison, for RBR and ROAR , we performed a hyperparameter search to find the most promising settings. In the case of RobX, we chose four combinations of the most relevant hyperparameters, $\tau$ and variance, where $\tau$ values were selected using the histogram technique described by the authors, and variances were chosen to vary significantly (specifically $0.1$ and $0.01$). The comprehensive description of the hyperparameter selection process is described in App. E.3. Note, that this comparison is not straightforward since each method is built on different premises and has multiple parameters to tune, making comparison to our probabilistic goals challenging.
We compare the above-listed baselines to two variants of our method ($\delta=0.8$, $\delta=0.9$) on the following quality measures. Empirical robustness, already defined in Eq. 8, quantifying the degree of robustness to model change. Proximity – the distance between the counterfactual and the original example:
$Proximity=d(x^{cf},x^{orig})$ (9)
In the experiments we calculate Proximity in two variants, $L_{1}$ and $L_{2}$, corresponding to Manhattan and Euclidean distance, respectively.
Plausibility – the average distance to $n$ closest neighbors of the CFE in the training set:
$Plausibility=\frac{1}{n}\sum_{i=1}^{n}d(x^{cf},x_{i})$ (10)
Distance to Base – a metric calculated only for the post-hoc methods, measuring the $L_{1}$ distance between the robust CFE and the base one:
$Distance~{}to~{}Base=d(x^{rcf},x^{bcf})$ (11)
Table 1: Comparative study results. RobX and BetaRCE were using GrowingSpheres as the base counterfactual explainer. Parameters used in a given method are listed next to this method’s name; for RobX these are $\tau$ and variance, while for BetaRCE $\delta$ (we omit $\alpha$ as it is fixed at $0.95$). The values in each cell represent the mean $\pm$ standard error. The column Type sorts the methods by categories. The abbreviations Btsr and Arch used next to BetaRCE in the Type column stand for Bootstrap and Architecture, respectively.
@@
BetaRCE Arch BetaRCE (0.8) 0.488 $\pm$ 0.013 0.870 $\pm$ 0.013 $0.382\pm 0.005$ 0.372 $\pm$ 0.004 0.966 $\pm$ 0.004 - -
BetaRCE (0.9) 0.607 $\pm$ 0.014 0.953 $\pm$ 0.013 $0.420\pm 0.006$ 0.378 $\pm$ 0.004 0.975 $\pm$ 0.004 - -
BetaRCE Btsr BetaRCE (0.8) 0.445 $\pm$ 0.006 0.840 $\pm$ 0.008 $0.359\pm 0.003$ 0.359 $\pm$ 0.002 - 0.903 $\pm$ 0.006 -
BetaRCE (0.9) 0.583 $\pm$ 0.007 0.949 $\pm$ 0.008 $0.407\pm 0.003$ 0.369 $\pm$ 0.002 - 0.937 $\pm$ 0.005 -
BetaRCE Seed BetaRCE (0.8) 0.247 $\pm$ 0.006 0.813 $\pm$ 0.008 $0.346\pm 0.003$ 0.350 $\pm$ 0.002 - - 0.871 $\pm$ 0.006
BetaRCE (0.9) 0.315 $\pm$ 0.006 0.862 $\pm$ 0.008 $0.367\pm 0.003$ 0.353 $\pm$ 0.002 - - 0.902 $\pm$ 0.006
@@
BetaRCE Arch BetaRCE (0.8) 1.538 $\pm$ 0.049 2.912 $\pm$ 0.053 $0.749\pm 0.014$ 0.802 $\pm$ 0.011 0.904 $\pm$ 0.006 - -
BetaRCE (0.9) 1.697 $\pm$ 0.031 2.927 $\pm$ 0.036 $0.753\pm 0.009$ 0.783 $\pm$ 0.007 0.935 $\pm$ 0.005 - -
BetaRCE Btsr BetaRCE (0.8) 2.288 $\pm$ 0.041 3.451 $\pm$ 0.044 $0.889\pm 0.011$ 0.859 $\pm$ 0.008 - 0.833 $\pm$ 0.007 -
BetaRCE (0.9) 3.547 $\pm$ 0.071 4.501 $\pm$ 0.073 $1.156\pm 0.019$ 1.044 $\pm$ 0.015 - 0.880 $\pm$ 0.006 -
BetaRCE Seed BetaRCE (0.8) 1.420 $\pm$ 0.021 2.526 $\pm$ 0.028 $0.653\pm 0.007$ 0.726 $\pm$ 0.004 - - 0.826 $\pm$ 0.007
BetaRCE (0.9) 1.927 $\pm$ 0.030 2.906 $\pm$ 0.035 $0.750\pm 0.009$ 0.776 $\pm$ 0.006 - - 0.902 $\pm$ 0.006
@@
The detailed results for Diabetes and HELOC datasets are presented in Tab. 1 and the results for the remaining datasets are in App. K. The results indicate that BetaRCE outperforms RobX in terms of the Distance to Base metric. Moreover, both the Proximity and Plausibility metrics are either preserved or slightly improved, validating our goal of not deteriorating the properties of the base counterfactual. This suggests that BetaRCE is effective at finding CFEs with a user-defined robustness level that are not significantly different from the base counterfactual, preserving its original properties and aligning with one of our primary motivations.
@@
Regarding Empirical Robustness, we observe that RobX often achieves a higher score than BetaRCE. However, this comes at a cost, as indicated by the fact that the Distance to Base metric for RobX is frequently worse than for BetaRCE. In some cases, RobX with well-selected hyperparameters achieves nearly perfect robustness but significantly deteriorates all other metrics (e.g., more than doubles the Distance to Base in comparison to BetaRCE ). Recall that the goal of BetaRCE is to have robustness greater than a specified lower bound while preserving as much of the base counterfactual as possible, and these goals have been achieved. The only setting where the lower bound is violated by the computed Empirical Robustness is the Bootstrap-HELOC combination for BetaRCE at $\delta=0.9$, where that value is slightly lower than the expected lower bound. However, the difference is negligible, and we attribute it to BetaRCE’s confidence level parameter.
@@
We also conducted additional experiments to verify the robustness of BetaRCE to a misdefined space of admissible model changes. The details are presented in App. F.4. In summary, although the theoretical probabilistic bounds do not hold for out-of-distribution model changes, they are often practically satisfied by BetaRCE . The results reveal good transferability of CFEs generated for Architecture and Bootstrap scenarios to Seed, as well as Seed CFEs to Architecture. The transfer to the Bootstrap scenario seems to be more difficult, sometimes resulting in a drop of Empirical Robustness as high as 4 p.p.
@@ 5 Final Remarks
The experiments have confirmed that the introduced probabilistic bounds, estimated in practice with a set of models, hold for the Empirical Robustness, as demonstrated across three types of model changes. Notably, counterfactuals generated by BetaRCE not only improved robustness, but also were closer to the base CFEs, retaining their properties better than CFEs generated by existing methods, while achieving a similar level of robustness.
Several avenues for future research emerge from our study. First, our approach relies on a fairly simple bootstrap estimation method to model the second-order probability distribution of robustness. While our experiments demonstrated the utility of this approach, exploring more sophisticated estimation methods could yield interesting insights. Additionally, the question of precisely quantifying model changes remains open. Specifically, distinguishing between slight, moderate, and substantial model changes – and determining appropriate responses to each – requires further investigation. Defining the space of admissible models more rigorously could enhance the precision of robustness estimates. Another potential direction is the use of auxiliary models to predict beta distribution parameters, which might replace querying a set of estimators, consequently accelerating optimization. Finally, examining alternative optimization algorithms could provide further opportunities to enhance the efficiency of BetaRCE.
Acknowledgments
This work was partially supported by the National Science Centre, Poland (Grant No. 2023/51/B/ST6/00545) and by Poznan University of Technology (internal grant no. SBAD/0752).
@@ Appendix A Reproducibility
To access the full version of the appendix, navigate to: https://ignacystepka.com/projects/betarce.html
A.1 Code Availability
To ensure reproducibility and enable further experimentation with BetaRCE, we make the source code publicly available on GitHub: https://github.com/istepka/betarce.
A.2 Datasets
Wine, Breast Cancer, Car eval, and Rice datasets were obtained from https://archive.ics.uci.edu/. Diabetes dataset was sourced from https://www.kaggle.com/datasets/mathchi/diabetes-data-set. HELOC (fico) dataset from https://community.fico.com/s/explainable-machine-learning-challenge?tabset-158d9=d157e.